OpenStudy (anonymous):

Find gof if f(x) = x + 7 and g(x) = 3x - 2.

OpenStudy (anonymous):

its is pretty simple gof = g(f(x)) g(f(x)) = g(x+7) = 3(x+7) -2 = 3x + 21 - 2 = 3x + 19

OpenStudy (anonymous):

Thank you :D your a life saver, i suck at math :p

OpenStudy (anonymous):

x) , you should learn some basics , its not that hard ..